OpenStudy (anonymous):

find the value of A in tan(A - 45 ) =1/3

OpenStudy (anonymous):

\[\tan (A-B)=\frac{\tan A-\tan B}{1+\tan A \tan B}\]

OpenStudy (anonymous):

@John4m2 now what will u get with that formula

OpenStudy (anonymous):

tanA=2

OpenStudy (anonymous):

sure i will get it thanks

OpenStudy (anonymous):

hey ademm it`s the value of a not tanA

OpenStudy (anonymous):

A=Arctan 2 use a calculator

OpenStudy (anonymous):

ok Mukushia thanks alot so it will be A =1/tan(2)